OpenStudy (anonymous):

Solve for the variable in this equation. x/2+1/4=3/4 x=

OpenStudy (anonymous):

l.c.d. = 4 2x/4=1/4+3/4=1 2x=4 x=2

OpenStudy (anonymous):

Thanks so much